OpenStudy (anonymous):

dy/dx = √ (xy+1). Would the answer be y/2√(xy+1)?

OpenStudy (anonymous):

\[\frac{ y }{ 2\sqrt{xy+1} }\]

ganeshie8 (ganeshie8):

is that a differential equation ? or u just wanto find derivative of : \(y = \sqrt{xy+1}\) ?

OpenStudy (anonymous):

implicit differentiation

OpenStudy (anonymous):

I think just needs the derivative but you're missing a -x in the numerator y/2square root(xy+1)-x

OpenStudy (anonymous):

denominator*

ganeshie8 (ganeshie8):

Alright, whats the actual input equation ? \(y = \sqrt{xy+1}\) or \(\frac{dy}{dx} = \sqrt{xy+1}\) ?

OpenStudy (anonymous):

\[y'(x)= \frac{ y }{ 2\sqrt{xy+1}-x }\]

ganeshie8 (ganeshie8):

if it is second one, you need to find second derivative i guess... could you pls clarify ha

OpenStudy (anonymous):

I'm so confused now >.<

OpenStudy (anonymous):

what??..sorru.

OpenStudy (anonymous):

Is your original question y = square root (xy+1) ? or is dy/dx

OpenStudy (anonymous):

dy/dx....

OpenStudy (anonymous):

wait no,it's said x= square root (xy+1)

OpenStudy (anonymous):

and then find dy/dx

OpenStudy (anonymous):

Then your answer is correct.

ganeshie8 (ganeshie8):

\(\large x = \sqrt{xy+1}\) differentiate bobth sides with.respect.to x \(\large 1 = \frac{1}{2\sqrt{xy+1}} (xy+1)'\)

ganeshie8 (ganeshie8):

^^that comes by chain rule differentiate the thing inside parenthesis, you wud get some dy/dx term, solve for it

OpenStudy (anonymous):

how would u do that?..

OpenStudy (anonymous):

Product rule

OpenStudy (anonymous):

nevermine i got it. Thanks!

ganeshie8 (ganeshie8):

\(\large x = \sqrt{xy+1}\) differentiate bobth sides with.respect.to x \(\large 1 = \frac{1}{2\sqrt{xy+1}} (xy+1)'\) \(\large 1 = \frac{1}{2\sqrt{xy+1}} \left((xy)' + (1)'\right) \) \(\large 1 = \frac{1}{2\sqrt{xy+1}} \left((xy)' + 0\right) \) \(\large 1 = \frac{1}{2\sqrt{xy+1}} \left((xy)' \right) \)

ganeshie8 (ganeshie8):

apply product rule for \((xy)'\)